OpenStudy (anonymous):

Evaluate, if possible: -√36

OpenStudy (anonymous):

-6

jimthompson5910 (jim_thompson5910):

\[-\sqrt{36}=-\sqrt{6^2}=-6\] So \[-\sqrt{36}=-6\]